OpenStudy (anonymous):

Solve these proportions: 3/x=x/4

OpenStudy (anonymous):

3=x^2/4 12=x^2 x=sqrt(12), x=-sqrt(12)

OpenStudy (anonymous):

the answer has to fit into a box that looks like this

OpenStudy (anonymous):

Hmm. I hate electronic mathematics, it just doesn't work. Try 1 sqrt( 12)

OpenStudy (anonymous):

see??

OpenStudy (anonymous):

do 2 sqrt 3

OpenStudy (anonymous):

Okie at the end of the lesson I'll let you know if it was correct or not

OpenStudy (anonymous):

Kk <3

OpenStudy (anonymous):

It was corrects ♥